# Sequential Conditional Transport on Probabilistic Graphs

### v2

@@ 5.1.1 General Context
Following Dwork et al. (2012), a fair decision means that “similar individuals" are treated similarly. As discussed in the introduction, Kusner et al. (2017) and Russell et al. (2017) considered a “counterfactual fairness" criterion. Based on the approach discussed above, it is possible to quantify unfairness, for a single individual, of a model $m$, trained on features $(s,\boldsymbol{x})$ to predict an outcome $y$. If $y\in\{0,1\}$ is binary, then $m$ represents the underlying score, corresponding to the conditional probability that $y=1$.
@@ 5.1.2 Illustration With Simulated Data
Iso-scores can be visualized at the top of Figure 7, with group 00 on the left, $1$ on the right. Consider an individual $(s,x_{1},x_{2})=(s=0,-2,-1)$ in group 00, with a score of $18.24\%$ (bottom left of Figure 7). Using Algorithm 1, its counterfactual counterpart $(s=1,x^{\star}_{1},x^{\star}_{2})$ can be constructed. The resulting score varies depending on the causal assumption. The score would be $40.95\%$ assuming the causal graph of Figure 5a, and $54.06\%$ assuming causal graph 5b. In the first case, the mutatis mutandis difference $m(s=1,x^{\star}_{1},x^{\star}_{2})-m(s=0,x_{1},x_{2})$, i.e., $+22.70\%$, is:
$m(s=1,x_{1},x_{2})-m(s=0,x_{1},x_{2})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ :-10.65\%$
$+$ $m(s=1,x^{\star}_{1},x_{2})-m(s=1,x_{1},x_{2})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ :+17.99\%$
$+$ $m(s=1,x^{\star}_{1},x^{\star}_{2})-m(s=1,x^{\star}_{1},x_{2})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ :+15.37\%.$
The first term is the ceteris paribus difference, the second one the change in $x_{1}$ and the third one the change in $x_{2}$, conditional on the change in $x_{1}$. If, instead, we assume the causal graph of Figure 5b, the score of the same individual would become $54.06\%$ and the mutatis mutandis difference $m(s=1,x^{\star}_{1},x^{\star}_{2})-m(s=0,x_{1},x_{2})$, i.e., $+35.82\%$, is:
@@
$+$ $m(s=1,x_{1},x_{2}^{\star})-m(s=1,x_{1},x_{2})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ :+16.07\%$
$+$ $m(s=1,x^{\star}_{1},x^{\star}_{2})-m(s=1,x_{1},x_{2}^{\star})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ :+30.41\%.$
@@ 5.2 Global Fairness Metrics
Instead of focusing on a single individual, it is possible to quantify the fairness of a model $m$ on a global scale. For example, the Demographic Parity criterion can be extended to Counterfactual Demographic Parity (CDP), allowing fairness assessment within a population subgroup with $s=0$. Consider the empirical version of “counterfactual fairness" in Kusner et al. (2017)
@@
which corresponds to the “average treatment effect of the treated" in the classical causal literature. This can be computed more efficiently using Algorithm 2 in Appendix B, which offers a faster alternative compared to Algorithm 1. Other group fairness metrics, based on Equalized Odds, can be extended to aggregated counterfactual fairness measures (see Appendix C).
@@ Conclusion
In this paper, we propose a sequential transport approach for constructing counterfactuals based on OT theory while respecting the underlying causal graph of the data. By using conditional univariate transport maps, we derive closed-form solutions for each coordinate of an individual’s characteristics, which facilitates the interpretation of both individual counterfactual fairness of our predictive model, and global fairness through “counterfactual demographic parity." Future work could extend counterfactual fairness evaluation to mitigation by applying pre-processing or in-processing methods using sequential transport for counterfactual generation.
@@ A.3 Conditional Gaussian Transport
Figure 10: Two Gaussian conditional OTs. On the left-hand side, the process begins with a univariate transport along the $x$ axis (using $T^{\star}_{x}$), followed by a transport along the $y$ axis on the conditional distributions (using $T_{y|x}^{\star}$), corresponding to the “lower triangular affine mapping." On the right-hand side, the sequence is reversed: it starts with a univariate transport along the $y$ axis (using $T^{\star}_{y}$) followed by transport along the $x$ axis on the conditional distributions (using $T_{x|y}^{\star}$). The red square is the multivariate OT of the point in the bottom left, corresponding to the “upper triangular affine mapping."

### v3

@@ 5.1.1 General Context
Following Dwork et al. (2012), a fair decision means that “similar individuals” are treated similarly. As discussed in the introduction, Kusner et al. (2017) and Russell et al. (2017) considered a “counterfactual fairness” criterion. Based on the approach discussed above, it is possible to quantify unfairness, for a single individual, of a model $m$, trained on features $(s,\boldsymbol{x})$ to predict an outcome $y$. If $y\in\{0,1\}$ is binary, then $m$ represents the underlying score, corresponding to the conditional probability that $y=1$.
@@ 5.1.2 Illustration With Simulated Data
Iso-scores can be visualized at the top of Figure 7, with group 00 on the left, $1$ on the right. Consider an individual $(s,x_{1},x_{2})=(s=0,-2,-1)$ in group 00, with a score of $18.24\%$ (bottom left of Figure 7). Using Algorithm 1, its counterfactual counterpart $(s=1,x^{\star}_{1},x^{\star}_{2})$ can be constructed. The resulting score varies depending on the causal assumption. The score would be $61.40\%$ assuming the causal graph of Figure 5a, and $56.34\%$ assuming causal graph 5b. In the first case, the mutatis mutandis difference $m(s=1,x^{\star}_{1},x^{\star}_{2})-m(s=0,x_{1},x_{2})$, i.e., $+43.15\%$, is:
@@
$+$ $m(s=1,x^{\star}_{1},x_{2})-m(s=1,x_{1},x_{2})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ :+15.63\%$
$+$ $m(s=1,x^{\star}_{1},x^{\star}_{2})-m(s=1,x^{\star}_{1},x_{2})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ :+38.18\%.$
The first term is the ceteris paribus difference, the second one the change in $x_{1}$ and the third one the change in $x_{2}$, conditional on the change in $x_{1}$. If, instead, we assume the causal graph of Figure 5b, the score of the same individual would become $56.34\%$ and the mutatis mutandis difference $m(s=1,x^{\star}_{1},x^{\star}_{2})-m(s=0,x_{1},x_{2})$, i.e., $+38.09\%$, is:
$+$ $m(s=1,x_{1},x_{2}^{\star})-m(s=1,x_{1},x_{2})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ :+14.51\%$
$+$ $m(s=1,x^{\star}_{1},x^{\star}_{2})-m(s=1,x_{1},x_{2}^{\star})\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ :+34.24\%.$
@@ 5.2 Global Fairness Metrics
Instead of focusing on a single individual, it is possible to quantify the fairness of a model $m$ on a global scale. For example, the Demographic Parity criterion can be extended to Counterfactual Demographic Parity (CDP), allowing fairness assessment within a population subgroup with $s=0$. Consider the empirical version of “counterfactual fairness” in Kusner et al. (2017)
@@
which corresponds to the “average treatment effect of the treated” in the classical causal literature. This can be computed more efficiently using Algorithm 2 in Appendix B, which offers a faster alternative compared to Algorithm 1. Other group fairness metrics, based on Equalized Odds, can be extended to aggregated counterfactual fairness measures (see Appendix C).
@@ Conclusion
In this paper, we propose a sequential transport approach for constructing counterfactuals based on OT theory while respecting the underlying causal graph of the data. By using conditional univariate transport maps, we derive closed-form solutions for each coordinate of an individual’s characteristics, which facilitates the interpretation of both individual counterfactual fairness of our predictive model, and global fairness through “counterfactual demographic parity.” Future work could extend counterfactual fairness evaluation to mitigation by applying pre-processing or in-processing methods using sequential transport for counterfactual generation.
@@ A.3 Conditional Gaussian Transport
Figure 10: Two Gaussian conditional OTs. On the left-hand side, the process begins with a univariate transport along the $x$ axis (using $T^{\star}_{x}$), followed by a transport along the $y$ axis on the conditional distributions (using $T_{y|x}^{\star}$), corresponding to the “lower triangular affine mapping.” On the right-hand side, the sequence is reversed: it starts with a univariate transport along the $y$ axis (using $T^{\star}_{y}$) followed by transport along the $x$ axis on the conditional distributions (using $T_{x|y}^{\star}$). The red square is the multivariate OT of the point in the bottom left, corresponding to the “upper triangular affine mapping.”